OpenStudy (anonymous):

A scientist uses a scale Y for measuring temperature. In this scale, pure water melts at 10oY and boils at 130oY at standard atmospheric pressure. The scientist measures the temperature at which sodium melts to be 127oY . Express this temperature in degrees Celsius 97.5oC 102.6oC 38.6oC 196.6oC

OpenStudy (anonymous):

If boiling temperature of water is 100C and in Yscale is 130Y then by ratio and proportion, \[1Y=100/130=\frac{ 10 }{ 13 }C\] For melting of sodium, the convertion should be \[127*\frac{ 10 }{ 13 }=97.7C\] Which is approximately 97.5C.

OpenStudy (anonymous):

@biafra2013 got it?